# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 7. Conclusion
This paper studies $\mathrm{RPLM}_{SED}$ relational prompt-based pre-trained language models for Social Event Detection. A pairwise message modeling strategy is proposed to address the issues of missing and noisy edges in social message graphs. A multi-relational prompt-based pairwise message learning mechanism is presented to simultaneously leverage the content and structural information of messages, learning more robust and stable message representations. A clustering constraint is designed to optimize the training process, enhancing message representations’ distinguishability. The conducted experiments on three real-world social media datasets demonstrate that $\mathrm{RPLM}_{SED}$ achieves SOTA performance in offline, online, low-resource, and long-tail distribution social event detection tasks. In the future, we aim to explore how $\mathrm{RPLM}_{SED}$ can be extended to semi-supervised or unsupervised social event detection tasks while also investigating ways to optimize the model architecture further to achieve more efficient detection.

### v2

@@ 7. Conclusion
This paper studies $\mathrm{RPLM}_{SED}$ relational prompt-based pre-trained language models for Social Event Detection. A pairwise message modeling strategy is proposed to address the issues of missing and noisy edges in social message graphs. A multi-relational prompt-based pairwise message learning mechanism is presented to simultaneously leverage the content and structural information of messages, learning more robust and stable message representations. A clustering constraint is designed to optimize the training process, enhancing message representations’ distinguishability. The conducted experiments on three real-world social media datasets demonstrate that $\mathrm{RPLM}_{SED}$ achieves SOTA performance in offline, online, low-resource, and long-tail distribution social event detection tasks. In the future, we aim to explore how supervised $\mathrm{RPLM}_{SED}$ can be extended to semi-supervised or unsupervised social event detection tasks while also investigating ways to optimize the model architecture further to achieve more efficient detection.
Acknowledgments
This work is supported by the NSFC through grants 62266028, 62322202, U21B2027, and 62432006, Yunnan Provincial Major Science and Technology Special Plan Projects through grants 202402AD080002, 202302AD080003 and 202202AD080003, Key Projects of Basic Research in Yunnan Province through grants 202301AS070047, 202301AT070471.